OpenStudy (anonymous):

how would i simplify c3+16c-2c2-32/-c2-16 using -1 as a factor

OpenStudy (anonymous):

-c^2 - 16 = -(c^2+16) then factorize the numerator usually there's a common factor between the numerator and denominator

OpenStudy (anonymous):

(c^2 +16)(c-2) / -(c^2+16) = -(c-2) = 2-c